# Rethinking HSM and TPM Security in the Cloud: Real-World Attacks and Next-Gen Defenses

### v1

@@ Abstract
As organizations rapidly migrate to the cloud, the security of cryptographic key management has become a growing concern. Hardware Security Modules (HSMs) and Trusted Platform Modules (TPMs), traditionally seen as the gold standard for securing encryption keys and digital trust, are increasingly challenged by cloud-native threats.
Real-world breaches have exposed weaknesses in cloud deployments, including misconfigurations, API abuse, and privilege escalations, allowing attackers to access sensitive key material and bypass protections. These incidents reveal that while the hardware remains secure, the surrounding cloud ecosystem introduces systemic vulnerabilities.
This paper analyzes notable security failures involving HSMs and TPMs, identifies common attack vectors, and questions longstanding assumptions about their effectiveness in distributed environments. We explore alternative approaches such as confidential computing, post-quantum cryptography, and decentralized key management.
Our findings highlight that while HSMs and TPMs still play a role, modern cloud security requires more adaptive, layered architectures. By evaluating both current weaknesses and emerging models, this research equips cloud architects and security engineers with strategies to reinforce cryptographic trust in the evolving threat landscape.
@@ 2.4 Industry-Wide Trends in Cloud Security Failures
Misconfigurations (IAM, firewall, storage) Bypasses key isolation and allows unauthorized access to encrypted data
API Exploits Allows remote actors to invoke sensitive operations such as key export or deletion
Privilege Escalation Grants administrative access over cryptographic modules and key material
Multi-Tenancy Risks Enables cross-tenant leakage or inference of protected cryptographic assets
@@ 3.1 HSMs: Strong Hardware, Soft Targets
Supply-Chain and API Exploits: A Wiz Security report highlights real-world cases where attackers exploited exposed CI / CD credentials and secrets, including those found in environment variables, build artifacts, or .bash_history files,to impersonate legitimate workloads and access HSM-backed interfaces [7]. These techniques bypass cryptographic enforcement not by breaking the HSM, but by abusing its trusted API surface.
@@
3.2 TPMs: Root-of-Trust, but Weak in the Cloud
@@
Core Strength Dedicated hardware for secure key management Hardware-rooted attestation, disk encryption, secure boot
Cloud Vulnerabilities API abuse, CI/CD token leakage, supply-chain compromise Hypervisor privilege, weak vTPM isolation, virtualization risks
@@
Operational Challenges Requires complex access control and monitoring Depends on TEE/hypervisor integrity; difficult to audit
Key Takeaway Hardware secure;but APIs and ops are weak links Theoretically sound;but cloud abstraction layers create attack surface
3.3 Conclusion: Ecosystem Failures, Not Hardware Flaws
Attackers do not need to break encryption or tamper with secure chips,instead, they exploit poorly scoped IAM permissions, insecure APIs, or compromised hypervisors. These indirect paths to key compromise render hardware protections ineffective unless paired with operational hardening, environment isolation, and consistent auditability.
@@ 4.5 Summary and Hybrid Approaches
Confidential Computing Isolates data-in-use with hardware Vulnerable to side-channels; TCB complexity
Post-Quantum Crypto Quantum-safe; NIST-backed Needs new hardware; bulky keys and compute costs
MPC/Decentralized Keys No single failure point; collaborative trust Operational complexity; coordination overhead
Secure vTPMs Boosts vTPM trust via enclaves Enclave overhead; setup pain

### v2

# HSM and TPM Failures in Cloud: A Real-World Taxonomy and Emerging Defenses

@@ Abstract
As cloud infrastructure becomes the backbone of modern organizations, the security of cryptographic key management, especially using Hardware Security Modules (HSMs) and Trusted Platform Modules (TPMs) faces unprecedented challenges. While these hardware-based solutions offer strong protection in isolated environments, their effectiveness is being undermined by cloud-native threats such as misconfigurations, compromised APIs, and lateral privilege escalations.
This paper presents a comprehensive analysis of publicly disclosed attacks and breaches involving HSMs and TPMs in cloud environments, identifying recurring architectural and operational flaws. We propose a taxonomy of attack vectors based on real-world case studies and threat intelligence reports, highlighting the gaps between hardware trust anchors and dynamic cloud ecosystems.
Furthermore, we evaluate emerging defensive paradigms: confidential computing, post-quantum cryptography, and decentralized key management systems (dKMS)—assessing their potential to address these gaps. Our findings emphasize that securing cloud-based cryptographic trust requires a layered, context-aware approach that integrates both hardware and software safeguards. The study serves as a practical framework for cloud architects and security engineers to reassess key protection strategies in light of evolving threats.
To our knowledge, this is the first work to synthesize documented, real-world cloud HSM and TPM failures into a coherent taxonomy grounded in modern threat models.
@@ 2.4 Industry-Wide Trends in Cloud Security Failures
Misconfigurations (IAM, firewall, storage) Can bypass key isolation, allowing unauthorized access to encrypted data.
API Exploits Remote attackers may invoke sensitive operations, such as key export or deletion.
Privilege Escalation Grants admin-level access to cryptographic modules and key material.
Multi-Tenancy Risks Potential for cross-tenant data leakage or inference of cryptographic assets.
@@ 3.1 HSMs: Strong Hardware, Soft Targets
Supply Chain and API Exploits: A Wiz Security report highlights real-world cases where attackers exploited exposed CI/CD credentials and secrets—found in environment variables, build artifacts, or .bash_history files—to impersonate legitimate workloads and access HSM-backed interfaces [7]. These techniques bypass cryptographic enforcement not by breaking the HSM, but by abusing its trusted API surface.
@@
3.2 TPMs: Root of Trust, but Weak in the Cloud
@@
3.3 The Real Problem: Ecosystem Failures
Attackers do not need to break encryption or tamper with secure chips. Instead, they exploit poorly scoped IAM permissions, insecure APIs, or compromised hypervisors. These indirect paths to key compromise render hardware protections ineffective unless paired with operational hardening, environment isolation, and consistent auditability.
@@ 3.3 The Real Problem: Ecosystem Failures
Core Strength Dedicated hardware for secure key management Hardware-rooted attestation, disk encryption, and secure boot
Cloud Vulnerabilities API abuse, CI/CD token leakage, supply chain compromise Hypervisor privilege escalation, weak vTPM isolation, virtualization risks
@@
Operational Challenges Requires complex access control and continuous monitoring Depends on TEE/hypervisor integrity; often difficult to audit
Key Takeaway Hardware-secure, but APIs and operational practices are weak links Theoretically sound, but cloud abstraction layers expand the attack surface
@@ 4.5 Summary and Hybrid Approaches
Confidential Computing Isolates data-in-use using trusted hardware enclaves Susceptible to side-channel attacks; large trusted computing base (TCB)
Post-Quantum Cryptography Resistant to quantum attacks; supported by NIST standardization Requires new algorithms, hardware support; larger keys and computational overhead
MPC / Decentralized Key Management Eliminates single point of failure; enables collaborative trust models High coordination overhead; complex implementation and maintenance
Secure vTPMs Strengthens vTPM trust using enclaves like Intel SGX or AMD SEV Enclave setup overhead; requires platform compatibility and tuning